# Local Level Dynamic Random Partition Models for Changepoint Detection

### v1

@@ 3 Posterior Inference
Posterior inference for the parameters of the LLDPM is carried out using a Markov chain Monte Carlo (MCMC) algorithm and, more specifically, a Gibbs sampling scheme (Neal, 2000). For computational convenience, we marginalize with respect to the local level parameters $\bm{\beta}_{t}$, thus introducing, for any $t=1,\ldots,T$, the conditional distribution of the observations $\bm{Y}_{t}$, given the partition $\pi_{t}$, namely
@@
The algorithm comprises three main steps, with particular attention needed for jointly updating the random partition $\pi_{t}$ and the changepoint $\gamma_{t}$ from their full conditional distribution. The updates also include the dependence parameters $\eta_{t}$, for which we specify a hyperprior distribution as $\eta_{t}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{Beta}(a,b)$.
We briefly describe the updates of the model parameters at a generic iteration. For simplicity, we do not introduce explicit notation to indicate at which iteration random quantities have been updated. Instead, the full conditional distributions for each update should be understood as being conditional on the most up-to-date values of the random quantities involved in the algorithm. The algorithm is presented for the case $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$. Extending it to other base partition distributions is conceptually straightforward. Additional details on the posterior distributions and our implementation can be found in Section B of the Supplementary Material.
@@
    Joint update of $(\pi_{t},\gamma_{t})$. For any $t=1,\ldots,T$, we follow a two-step procedure.
@@
Step 1.1a highlights the computational challenges that arise when implementing our PSM. Although the partition-based state equation (3) has the same mixture representation that characterizes tractable priors, such as the spike-and-slab prior commonly used for variable selection, working with random partitions makes computations more complex. The time-specific marginal likelihood (14) is estimated via Monte Carlo, a step that, conveniently, must be performed only once, prior to running the Gibbs sampler. Section B.1 in the Supplementary Material provides further details on how this step can be made efficient. Step 1.1a also involves sampling from a distribution proportional to $p_{\text{CRP}}(\pi_{t})p(\bm{Y}_{t}\mid\pi_{t})$, for every $t=1,\ldots,T$. We propose implementing an auxiliary MCMC run prior to initializing the primary MCMC algorithm for model fitting. In this preliminary run, we fit an independent CRP model at each time point. We then save all the partitions generated throughout the MCMC iterations, creating, for each $t=1,\ldots,T$, a catalogue $\mathscr{S}_{t}$ of realizations from the posterior distribution of $\pi_{t}$ given $\bm{Y}_{t}$. The size of these catalogues can be arbitrarily large. To sample a new partition from a distribution proportional to $p_{\text{CRP}}(\pi_{t})p(\bm{Y}_{t}\mid\pi_{t})$, we refer to the relevant catalogue of realized partitions from the auxiliary MCMC run, randomly selecting a new partition from $\mathscr{S}_{t}$, for $t=1,\ldots,T$. This strategy is crucial for making the resulting MCMC algorithm efficient and capable of handling observations recorded over a large number of time points $T$. For instance, this is the case in the application presented in Section 5, where $T=349$.
@@ 3.1 Changepoint detection
Changepoint detection inherently involves making multiple comparisons, since the decisions are temporally dependent. To address this multi-comparison problem, we use a compound decision-theoretic approach to detect the presence of a changepoint, which is based on a loss function that takes simultaneously into account the sequence of decisions, and it is defined as a linear combination of measures of the false positive and true positive (or, alternatively, false negative) decisions (Sun and Cai, 2007). In a Bayesian context, Müller et al. (2004) and Müller et al. (2006) have demonstrated that, under the assumption of both independent hypotheses and independent (marginal) loss functions, the optimal approach for minimizing the resulting posterior expected loss involves thresholding the posterior probabilities of the changepoint ($\text{PPC}_{t}$), $\text{Pr}(\gamma_{t}=1\mid\bm{Y}_{1},\ldots,\bm{Y}_{T})$, for $t=1,\ldots,T$, as estimated from the MCMC output. However, such a procedure does not inherently control for a given false discovery rate (FDR) unless such control is explicitly accounted for. That is, we need to determine the optimal threshold in order to control the FDR at a specific desired level, say $\zeta$. More in detail, we consider the Bayesian FDR (Newton et al., 2004),
$\text{BFDR}_{\text{m}}(h)=\frac{\sum_{t=1}^{T}(1-\text{PPC}_{t})\mathbbm{1}_{\{\text{PPC}_{t}>h\}}}{\max\left(\sum_{t=1}^{T}\mathbbm{1}_{\{\text{PPC}_{t}>h\}},1\right)},$ (15)
@@
where $h$ is the chosen threshold and $\mathbbm{1}_{\{\text{PPC}_{t}>h\}}$ is an indicator function such that $\mathbbm{1}_{\{\text{PPC}_{t}>h\}}=1$ if $\text{PPC}_{t}>h$, and 0 otherwise. The optimal threshold $h^{*}$ corresponds to the minimum value of $h$ that ensures that the $\mathrm{BFDR}_{\mathrm{m}}$ is less than $\zeta$. In formulas, $h^{*}=\min\{h:\operatorname{BFDR}_{\text{m}}(h)\leq\zeta\}$. The previous testing procedure can be classified as a marginal approach (hence, the subscript in $\text{BFDR}_{\text{m}}$) since it fails to consider existing dependences either among hypotheses or in the decisions themselves. Sun et al. (2015) extended this framework to the spatial setting, explicitly taking into account dependences among the hypotheses, as induced by a spatial model. More recently, Chandra and Bhattacharya (2019) introduced non-marginal loss functions and non-marginal decision rules that directly account for dependences in the decision-making process. Their procedure incorporates additional information about dependences among tests into the definition of error and non-error terms for subgroups of hypotheses. Specifically, the approach penalizes each hypothesis based on incorrect decisions related to other dependent tests, thus defining a compound loss where decisions about dependent tests are interrelated. We adapt their framework to our case. For each time point $t=1,\ldots,T$, the detection of a changepoint at $t$ is related to the probability of a changepoint at times $t-1$ and $t+1$. That is, the null hypothesis $H_{0,t}$, i.e. no changepoint at time $t$, is assessed with respect to the related decisions at times $t-1$ and $t+1$. Considering this set of dependent hypotheses together is crucial, since, for example, a false changepoint detection at time $t-1$ may induce a false changepoint detection at time $t$, even if the null hypothesis is true at both times. We let $d_{t}$ represent the decision at time $t$, i.e., $d_{t}=1$ if the $t$-th hypothesis is rejected and $d_{t}=0$ if it is not. Similarly, $r_{t}$ denotes the truth at time $t$, i.e., $r_{t}=1$ if $H_{0,t}$ is true, $r_{t}=0$ otherwise. We also introduce the notation $\bm{d}=(d_{1},\ldots,d_{T})$, $\bm{r}=(r_{1},\ldots,r_{T})$. We define the true positive rate (TPR) as the ratio between the number of times where the decision correctly identifies a changepoint, and the number of positive decisions. That is
$\text{TPR}(\bm{d},\bm{r})=\frac{1}{D}\sum_{t=1}^{T}d_{t}r_{t},$
where $D=\sum_{t=1}^{T}d_{t}$. We consider the compound loss function
$L(\bm{d},\bm{r})=-\text{TPR}(\bm{d},\bm{r})+\lambda\text{ER}(\bm{d},\bm{r}),$
@@
where $\lambda$ is a positive constant, and, in order to penalize false detections at each time $t$, the error rate (ER) is defined as the ratio between the total number of false detections in the set of timepoints $\mathcal{G}_{t}$ and the number of positive decisions $D$. That is
$\text{ER}(\bm{d},\bm{r})$ $=\frac{1}{D}\left\{\sum_{t=2}^{T}d_{t-1}(1-r_{t-1})+\sum_{t=1}^{T}d_{t}(1-r_{t})+\sum_{t=1}^{T-1}d_{t+1}(1-r_{t+1})\right\}$ (16)
$=\frac{1}{D}\left\{d_{1}(1-r_{1})+3\sum_{t=1}^{T}d_{t}(1-r_{t})+d_{T}(1-r_{T})\right\}.$
@@
We aim to minimize the posterior expected loss with respect to $\bm{d}$. Then, following steps similar to those in Theorem 1 of Müller et al. (2004), it is possible to show that the optimal decision rule coincides with setting a threshold on the posterior probabilities $\text{PPC}_{t}$. Additionally, because the expression for the error rate in equation (16) includes the term $3\sum_{t=1}^{T}d_{t}(1-r_{t})$, the resulting non-marginal Bayesian FDR, denoted as $\text{BFDR}_{\text{nm}}$, can be controlled such that $\text{BFDR}_{\text{nm}}(h)=3\,\text{BFDR}_{\text{m}}(h)$, where $\text{BFDR}_{\text{m}}$ is defined in (15). Thus, it is possible to control the non-marginal $\text{BFDR}_{\text{nm}}$ by computing the commonly used marginal Bayesian FDR and considering a more stringent level, say $\zeta/3$, for the identification of the optimal threshold on the $\text{PPC}_{t}$’s. This correction becomes especially relevant when dealing with autoregressive data, which are characterized by higher structural dependence.
@@ 4 Simulation studies
We present two simulation studies to illustrate the performance of our LLDPM when analyzing data from different data-generating mechanisms. Specifically, we investigate the model’s ability to accurately detect changepoints in both independent and autocorrelated data scenarios, while also recovering the time-specific latent cluster structure. For comparison, we evaluate our model against four alternatives:
@@
    The Linear Dependent Dirichlet Process (LDDP) by Quintana et al. (2022), which incorporates time explicitly within the atoms of the dependent process.
@@
It is important to note that, among the four alternatives we consider, the last three models are not explicitly designed to model the evolution of random partitions. To implement them for the purpose of changepoint detection, the data must be consolidated into a vector with time treated as a covariate. In contrast, the DRPM and our model are inherently model-based and do not require data concatenation to introduce time dependence. Additionally, since the alternative approaches are not specifically designed to detect changepoints in the evolution of partitions over time, they typically identify many changepoints (false detections), unless we adopt a more liberal approach when calculating the similarity between partitions at consecutive time steps. Specifically, we identify a changepoint only if the similarity between partitions at consecutive time points is less than 90% based on the ARI. This rule seemed to enhance the performance of all competitor models.
@@ 4.1 Simulations with independent data
$n$ Measure LLDPM DRPM LDDP WDDP GMDDP 20 specificity 0.6104 (0.413) 0.6106 (0.413) 0.566 (0.288) 0.3796 (0.021) 1.000 (0.000) accuracy 0.9902 (0.010) 0.6288 (0.375) 0.6008 (0.265) 0.4292 (0.019) 1.000 (0.000) recall 1.000 (0.000) 0.8375 (0.196) 1.000 (0.000) 1.000 (0.000) 1.000 (0.000) precision 0.9014 (0.094) 0.4209 (0.367) 0.2714 (0.238) 0.123 (0.004) 1.000 (0.000) F1 0.9455 (0.054) 0.4724 (0.337) 0.3836 (0.243) 0.2191 (0.006) 1.000 (0.000) AUC 0.9947 (0.006) 0.7263 (0.194) 0.6472 (0.176) 0.6898 (0.010) 1.000 (0.000) 50 specificity 0.6309 (0.396) 0.6317 (0.397) 0.1961 (0.131) 0.0233 (0.021) 0.9989 (0.008) accuracy 0.9664 (0.023) 0.6558 (0.363) 0.2604 (0.121) 0.1014 (0.020) 0.9990 (0.007) recall 1.000 (0.000) 0.9325 (0.098) 1.000 (0.000) 1.000 (0.000) 1.000 (0.000) precision 0.7331 (0.149) 0.4292 (0.343) 0.0995 (0.013) 0.0818 (0.002) 0.9924 (0.054) F1 0.8378 (0.097) 0.5123 (0.330) 0.1808 (0.022) 0.1512 (0.002) 0.9952 (0.034) AUC 0.9817 (0.012) 0.7821 (0.193) 0.5961 (0.066) 0.5116 (0.011) 0.9995 (0.004) 100 specificity 0.6228 (0.355) 0.6146 (0.352) 0.1397 (0.126) 0.02717 (0.022) 0.9839 (0.080) accuracy 0.9514 (0.031) 0.6390 (0.321) 0.2086 (0.116) 0.1050 (0.020) 0.9852 (0.073) recall 1.0000 (0.000) 0.9200 (0.123) 1.0000 (0.000) 1.0000 (0.000) 1.0000 (0.000) precision 0.6611 (0.167) 0.3444 (0.281) 0.0934 (0.013) 0.0821 (0.002) 0.9671 (0.163) F1 0.7843 (0.118) 0.4408 (0.281) 0.1707 (0.021) 0.1517 (0.003) 0.9721 (0.138) AUC 0.9740 (0.017) 0.7673 (0.168) 0.5685 (0.063) 0.5134 (0.011) 0.9920 (0.040)
@@
Table 1: Section 4.1. Performance measures for changepoint detection with independent data, for the LLDPM and four competing models. The values correspond to the average (standard errors) over 50 simulations.
@@
We begin by simulating independent data characterized by similar partition structures over different intervals of time, trying to mimic the data-generating process characterizing the LLDPM. We evaluate performance across 50 replicated datasets and different numbers of observations, specifically $n=\{20,50,100\}$, over $T=100$ time points. In each scenario, we consider data partitions characterized by eight changepoints. At each time point, given the time-specific partition, we generate the observations for each cluster from a Normal distribution. The means of these are simulated from a Normal distribution with zero mean and variance 0.25, while the variance is common across components and set to 0.01. An example of the generated data is shown in Figure C.8 in the Supplementary Material.

### v2

@@ 3 Posterior Inference
[…]
@@
The algorithm comprises three main steps, with particular attention needed for jointly updating the random partition $\pi_{t}$ and the changepoint $\gamma_{t}$ from their full conditional distribution. The updates also include the dependence parameters $\eta_{t}$, for which we specify a hyperprior distribution as $\eta_{t}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{Beta}(a,b)$. We briefly describe the updates of the model parameters at a generic iteration. At each iteration the random parameters involved in the MCMC scheme are the partitions $\bm{\pi}_{1:T}$, the auxiliary variable $\bm{\gamma}_{2:T}$, the dependence parameter $\eta_{2},\ldots,\eta_{T}$ and the local level parameters $\bm{\beta}_{1},\ldots,\bm{\beta}_{T}$. For simplicity, we do not introduce explicit notation to indicate at which iteration random quantities have been updated. Instead, the full conditional distributions for each update should be understood as being conditional on the most up-to-date values of the random quantities involved in the algorithm. The algorithm is presented for the case $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$. Extending it to other base partition distributions is conceptually straightforward. Additional details on the posterior distributions and our implementation can be found in Section B of the Supplementary Material.
@@
    Update of $(\pi_{t},\gamma_{t})$. For each $t=1,\ldots,T$, we follow a two-step procedure, detailed below for $t=2,\ldots,T-1$, with straightforward adjustments for the boundary cases $t=1$ and $t=T$.
@@
Step 1.1a highlights the computational challenges that arise when implementing our PSM. Although the partition-based state equation (3) has the same mixture representation that characterizes tractable priors, such as the spike-and-slab prior commonly used for variable selection, working with random partitions makes computations more complex. The time-specific marginal likelihood (16) is estimated via Monte Carlo, a step that, conveniently, must be performed only once, prior to running the Gibbs sampler. Section B.1 in the Supplementary Material provides further details on how this step can be made efficient. Step 1.1a also involves sampling from a distribution proportional to $p_{\text{CRP}}(\pi_{t})p(\bm{Y}_{t}\mid\pi_{t})$, for every $t=1,\ldots,T$. We propose implementing an auxiliary MCMC run prior to initializing the primary MCMC algorithm for model fitting. In this preliminary run, we fit an independent CRP model at each time point. We then save all the partitions generated throughout the MCMC iterations, creating, for each $t=1,\ldots,T$, a catalogue $\mathscr{S}_{t}$ of realizations from the posterior distribution of $\pi_{t}$ given $\bm{Y}_{t}$. The size of these catalogues can be arbitrarily large. To sample a new partition from a distribution proportional to $p_{\text{CRP}}(\pi_{t})p(\bm{Y}_{t}\mid\pi_{t})$, we refer to the relevant catalogue of realized partitions from the auxiliary MCMC run, randomly selecting a new partition from $\mathscr{S}_{t}$, for $t=1,\ldots,T$. This strategy is crucial for making the resulting MCMC algorithm efficient and capable of handling observations recorded over a large number of time points $T$. For instance, this is the case in the application presented in Section 5.1, where $T=349$. The described MCMC algorithm was implemented using the R software (R Core Team, 2021) combined with the C++ programming language via RcppArmadillo (Eddelbuettel and Sanderson, 2014). For a dataset consisting of $100$ time series observed over $100$ time points, the average computational time required for $1,000$ iterations is approximately $951.1$ seconds on a machine with an i7 processor and 16GB of RAM. Generalizing the presented algorithm to the multiview setting requires additional considerations. While it adapts naturally to two-view scenarios with minimal adjustments, as displayed in the analysis presented in Section 5.2, generalizing it to more than two views requires a more deliberate approach. This task involves identifying a common partition framework, which may serve as the basis for aligning and comparing different views. Thus, extending the method to settings with more than two views introduces additional computational complexity, which must be carefully managed to maintain the algorithm’s efficiency and robustness.
@@ 3.1 Changepoint detection
Changepoint detection inherently involves making multiple comparisons, since the decisions are temporally dependent. To address this multi-comparison problem, we use a compound decision-theoretic approach to detect the presence of a changepoint, which is based on a loss function that takes simultaneously into account the sequence of decisions, and it is defined as a linear combination of measures of the false positive and true positive (or, alternatively, false negative) decisions (Sun and Cai, 2007). In a Bayesian context, Müller et al. (2004) and Müller et al. (2006) have demonstrated that, under the assumption of both independent hypotheses and independent (marginal) loss functions, the optimal approach for minimizing the resulting posterior expected loss involves thresholding the posterior probabilities of the changepoint ($\text{PPC}_{t}$), $\text{Pr}(\gamma_{t}=1\mid\bm{Y}_{1:T})$, for $t=2,\ldots,T$, as estimated from the MCMC output. However, such a procedure does not inherently control for a given false discovery rate (FDR) unless such control is explicitly accounted for. That is, we need to determine the optimal threshold in order to control the FDR at a specific desired level, say $\zeta$. More in detail, we consider the Bayesian FDR (Newton et al., 2004),
$\text{BFDR}_{\text{m}}(h)=\frac{\sum_{t=2}^{T}(1-\text{PPC}_{t})\mathbbm{1}_{\{\text{PPC}_{t}>h\}}}{\max\left(\sum_{t=2}^{T}\mathbbm{1}_{\{\text{PPC}_{t}>h\}},1\right)},$ (17)
@@
where $h$ is the chosen threshold and $\mathbbm{1}_{\{\text{PPC}_{t}>h\}}$ is an indicator function such that $\mathbbm{1}_{\{\text{PPC}_{t}>h\}}=1$ if $\text{PPC}_{t}>h$, and 0 otherwise. The optimal threshold $h^{*}$ corresponds to the minimum value of $h$ that ensures that the $\mathrm{BFDR}_{\mathrm{m}}$ is less than $\zeta$. In formulas, $h^{*}=\min\{h:\operatorname{BFDR}_{\text{m}}(h)\leq\zeta\}$. The previous testing procedure can be classified as a marginal approach (hence, the subscript in $\text{BFDR}_{\text{m}}$) since it fails to consider existing dependences either among hypotheses or in the decisions themselves. Sun et al. (2015) extended this framework to the spatial setting, explicitly taking into account dependences among the hypotheses, as induced by a spatial model. More recently, Chandra and Bhattacharya (2019) introduced non-marginal loss functions and non-marginal decision rules that directly account for dependences in the decision-making process. Their procedure incorporates additional information about dependences among tests into the definition of error and non-error terms for subgroups of hypotheses. Specifically, the approach penalizes each hypothesis based on incorrect decisions related to other dependent tests, thus defining a compound loss where decisions about dependent tests are interrelated. We adapt their framework to our case. For each time point $t=2,\ldots,T$, the detection of a changepoint at $t$ is related to the probability of a changepoint at times $t-1$ and $t+1$. That is, the null hypothesis $H_{0,t}$, i.e. no changepoint at time $t$, is assessed with respect to the related decisions at times $t-1$ and $t+1$. Considering this set of dependent hypotheses together is crucial, since, for example, a false changepoint detection at time $t-1$ may induce a false changepoint detection at time $t$, even if the null hypothesis is true at both times. We let $d_{t}$ represent the decision at time $t$, i.e., $d_{t}=1$ if the $t$-th hypothesis is rejected and $d_{t}=0$ if it is not. Similarly, $r_{t}$ denotes the truth at time $t$, i.e., $r_{t}=0$ if $H_{0,t}$ is true, $r_{t}=1$ otherwise. We also introduce the notation $\bm{d}=(d_{2},\ldots,d_{T})$, $\bm{r}=(r_{2},\ldots,r_{T})$. We define the true positive rate (TPR) as the ratio between the number of times where the decision correctly identifies a changepoint, and the number of positive decisions. That is
$\text{TPR}(\bm{d},\bm{r})=\frac{1}{D}\sum_{t=2}^{T}d_{t}r_{t},$
where $D=\sum_{t=2}^{T}d_{t}$. We consider the compound loss function
$L(\bm{d},\bm{r})=-\text{TPR}(\bm{d},\bm{r})+\kappa\,\text{ER}(\bm{d},\bm{r}),$
@@
where $\kappa$ is a positive constant, and, in order to penalize false detections at each time $t$, the error rate (ER) is defined as the ratio between the total number of false detections in the set of time points $\mathcal{G}_{t}$ and the number of positive decisions $D$. That is,
$\text{ER}(\bm{d},\bm{r})$ $=\frac{1}{D}\left\{\sum_{t=3}^{T}d_{t-1}(1-r_{t-1})+\sum_{t=2}^{T}d_{t}(1-r_{t})+\sum_{t=2}^{T-1}d_{t+1}(1-r_{t+1})\right\}$ (18)
$=\frac{1}{D}\left\{d_{2}(1-r_{2})+3\sum_{t=2}^{T}d_{t}(1-r_{t})+d_{T}(1-r_{T})\right\}.$
While the true changepoint times $t$, where $r_{t}=1$, are unknown, they can be estimated using the corresponding changepoint indicators $\gamma_{t}$. Therefore, we seek to minimize the posterior expected loss $\mathbb{E}[L(\bm{d},\bm{\gamma})\mid\bm{Y}_{1:T}]$ with respect to $\bm{d}$, where the expectation is taken over the posterior distribution of $\bm{\gamma}$, which is compact notation for $\bm{\gamma}_{2:T}$. This leads to the following optimization problem for identifying the changepoints:
$\min_{\bm{d}}\,\mathbb{E}[L(\bm{d},\bm{\gamma})\mid\bm{Y}_{1:T}]$ $=\min_{\bm{d}}\left\{-\mathbb{E}[\text{TPR}(\bm{d},\bm{\gamma})\mid\bm{Y}_{1:T}]+\kappa\,\mathbb{E}[\text{ER}(\bm{d},\bm{\gamma}_{2:T})\mid\bm{Y}_{1:T}]\right\}$
$=\min_{\bm{d}}\left\{-\text{TPR}(\bm{d},\bm{\text{PPC}})+\kappa\,\text{ER}(\bm{d},\bm{\text{PPC}})\right\},$
@@
where PPC is the vector of posterior probabilities of a changepoint, $\text{PPC}_{t}$, $t=2,\ldots,T$. Following a similar approach to Theorem 1 in Müller et al. (2004), it can be shown that the optimal decision rule corresponds to setting a threshold on the posterior probabilities $\text{PPC}_{t}$. Additionally, because the expression for the error rate in equation (18) includes the term $3\sum_{t=2}^{T}d_{t}(1-r_{t})$, the resulting non-marginal Bayesian FDR, denoted as $\text{BFDR}_{\text{nm}}$, can be controlled such that $\text{BFDR}_{\text{nm}}(h)=3\,\text{BFDR}_{\text{m}}(h)$, where $\text{BFDR}_{\text{m}}$ is defined in (17). Thus, it is possible to control the non-marginal $\text{BFDR}_{\text{nm}}$ by computing the commonly used marginal Bayesian FDR and considering a more stringent level, say $\zeta/3$, for the identification of the optimal threshold on the $\text{PPC}_{t}$’s. This correction becomes especially relevant when dealing with autoregressive data, which are characterized by higher structural dependence.
@@ 4 Simulation studies
We present two simulation studies to illustrate the performance of our LLDPM under different data-generating mechanisms.We consider two versions of the LLDPM, defined by the choice of base partition distribution. In the first version, denoted LLDPM0, we set $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$, corresponding to a CRP. In the second version, denoted LLDPM0.25, we use a 2-CRP and set $p^{*}(\cdot)=p_{\sigma}(\cdot)$ with a discount parameter $\sigma=0.25$. Such choice of $\sigma$ is arbitrary and simply intended to demonstrate the applicability of our method when the base partition distribution differs from the CRP. While other choices are possible, additional investigations, not reported here, suggest that the method is rather robust to such variations. We then investigate the models ability to accurately detect changepoints in both independent and autocorrelated data scenarios, while also recovering the time-specific latent cluster structure. For comparison, we evaluate our models against six alternatives:
@@
    The Linear Dependent Dirichlet Process (LDDP) described by De Iorio et al. (2004), which incorporates time explicitly within the atoms of the dependent process.
@@
  5. 5.
    The Product Partition Model (PPM) implemented in the R package ppmSuite (Page and Quinlan, 2022), a Bayesian model where the set of changepoint indicators between time series are correlated.
  6. 6.
    The Autoregressive Dirichlet Process (AR1-DP) introduced by De Iorio et al. (2023), which adopts an autoregressive prior to allow for temporal dynamic clustering.
It is important to emphasize that among the six alternatives considered, only the DRPM is explicitly designed to model the evolution of random partitions, as is the case for our proposed model. The remaining models were not devised for this purpose. Specifically, the PPM is designed for multivariate time series and detects changepoints as shifts in the mean of the process rather than changes in partitions. The LDDP, WDDP, and GMDDP models are adapted for changepoint detection by organizing the data as a single vector and incorporating time as a covariate, following the approach of Page et al. (2022). The AR1-DP considers a sequence of autoregressive random probability measures over time, but does not explicitly incorporate changepoints detection. To identify changepoints when using models that are not specifically designed to detect changes in the partition evolution over time, we compare cluster configurations between consecutive time points: if the partitions are identical, no changepoint is recorded. However, applying this procedure often leads to an overidentification of changepoints. To address this limitation, we adopt a more flexible, similarity-based criterion for changepoint detection in the competitor models DRPM, LDDP, WDDP, GMDDP and AR1-DP. Specifically and only for these methods, a changepoint is identified when the ARI between consecutive partitions falls below 85%, in combination with the same FDR correction adopted for our LLDPM. This adjustment leads to improved performance for these methods (see Supplementary Material, Section C.1, for details). For the PPM model, we adopt a different strategy and rely on the estimated individual changepoints, as the model does not account for time-specific partitions.
We evaluate the performance of the above methods on two key tasks: accurately estimating the true partition of subjects at each time point and correctly identifying changepoints in partition evolution over time. For the first task, we use the ARI to compare the true partition with the one estimated by minimizing the lower bound of the posterior expected Variation of Information (VI; Wade and Ghahramani, 2018). For the second task, we use six standard metrics, namely specificity, accuracy, recall, precision, F1-score, and AUC, as defined in Section C.1 of the Supplementary Material. For the PPM model, these metrics are first computed at the individual changepoint level and then averaged across the entire sample. Finally, due to its autoregressive structure, the AR1-DP is applied exclusively in the autoregressive data scenario presented in Section 4.2.
@@ 4.1 Simulations with independent data
$n$ Measure LLDPM0 LLDPM0.25 DRPM LDDP WDDP GMDDP PPM 20 specificity 0.99 (0.01) 0.97 (0.03) 0.99 (0.01) 0.91 (0.17) 0.47 (0.04) $\mathbf{1.00}$ (0.00) 0.77 (0.01) accuracy 0.99 (0.01) 0.97 (0.03) 0.96 (0.01) 0.92 (0.16) 0.52 (0.03) $\mathbf{1.00}$ (0.00) 0.72 (0.01) recall $\mathbf{1.00}$ (0.00) $\mathbf{1.00}$ (0.00) 0.63 (0.14) $\mathbf{1.00}$ (0.00) 1.00 (0.02) $\mathbf{1.00}$ (0.00) 0.23 (0.01) precision 0.90 (0.09) 0.76 (0.16) 0.89 (0.12) 0.81 (0.32) 0.14 (0.01) $\mathbf{1.00}$ (0.00) 0.08 (0.00) F1 0.95 (0.05) 0.85 (0.11) 0.73 (0.12) 0.85 (0.28) 0.25 (0.01) $\mathbf{1.00}$ (0.00) 0.13 (0.01) AUC 0.99 (0.01) 0.98 (0.01) 0.80 (0.08) 0.85 (0.21) 0.73 (0.05) $\mathbf{1.00}$ (0.00) 0.50 (0.01) 50 specificity 0.96 (0.03) 0.91 (0.08) 0.99 (0.01) 0.44 (0.40) 0.29 (0.10) $\mathbf{1.00}$ (0.03) 0.66 (0.00) accuracy 0.97 (0.02) 0.91 (0.07) 0.96 (0.36) 0.49 (0.36) 0.35 (0.09) $\mathbf{1.00}$ (0.02) 0.63 (0.00) recall $\mathbf{1.00}$ (0.00) $\mathbf{1.00}$ (0.00) 0.67 (0.15) $\mathbf{1.00}$ (0.00) 1.00 (0.02) $\mathbf{1.00}$ (0.00) 0.34 (0.00) precision 0.73 (0.15) 0.56 (0.23) 0.88 (0.11) 0.34 (0.38) 0.11 (0.01) $\mathbf{0.99}$ (0.09) 0.08 (0.00) F1 0.84 (0.10) 0.70 (0.18) 0.74 (0.10) 0.41 (0.35) 0.20 (0.02) $\mathbf{0.99}$ (0.07) 0.14 (0.00) AUC 0.98 (0.01) 0.91 (0.14) 0.83 (0.07) 0.68 (0.20) 0.64 (0.05) $\mathbf{1.00}$ (0.01) 0.50 (0.00) 100 specificity 0.95 (0.03) 0.83 (0.10) $\mathbf{0.99}$ (0.01) 0.38 (0.36) 0.23 (0.04) 0.96 (0.11) 0.58 (0.00) accuracy 0.95 (0.03) 0.85 (0.09) $\mathbf{0.96}$ (0.02) 0.43 (0.33) 0.28 (0.04) 0.28 (0.10) 0.56 (0.00) recall $\mathbf{1.00}$ (0.00) $\mathbf{1.00}$ (0.00) 0.60 (0.15) $\mathbf{1.00}$ (0.00) 0.88 (0.03) $\mathbf{1.00}$ (0.00) 0.44 (0.00) precision 0.66 (0.17) 0.40 (0.16) 0.81 (0.15) 0.26 (0.32) 0.09 (0.01) $\mathbf{0.92}$ (0.24) 0.08 (0.00) F1 0.78 (0.12) 0.55 (0.16) 0.68 (0.13) 0.34 (0.30) 0.16 (0.01) $\mathbf{0.93}$ (0.20) 0.14 (0.00) AUC 0.97 (0.02) 0.83 (0.17) 0.79 (0.08) 0.67 (0.18) 0.55 (0.02) $\mathbf{0.98}$ (0.06) 0.52 (0.00)
@@
Table 1: Section 4.1. Performance measures for changepoint detection with independent data, for the LLDPM under two specifications and five competing models. The values correspond to the average (standard deviation) over 50 replicates.
@@
We begin by simulating independent data characterized by similar partition structures over different intervals of time, trying to mimic the data-generating process characterizing the LLDPM. We evaluate performance across 50 replicated datasets and different numbers of observations, specifically $n=\{20,50,100\}$, over $T=100$ time points. In each scenario, we consider data partitions characterized by eight changepoints. The time axis is divided into nine consecutive blocks of varying lengths, each representing a segment of the 100 total time points. Within each block, the partition of the units remains fixed and corresponds to one of three configurations. Two of these configurations consist of three clusters and differ only in the assignment of units to clusters, while the third configuration includes just two clusters. These configurations alternate across blocks. At each time point, given the time-specific partition defined by the active configuration, observations for each cluster are drawn from a Normal distribution. The variance is fixed at 0.01 and shared across all clusters, while the mean for each cluster is independently sampled from a Normal distribution with arbitrarily chosen parameters. An example of the generated data is shown in Figure C.9 in the Supplementary Material.